# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 3. Overview of LLM Vulnerabilities
In recent studies, the vulnerabilities and challenges of LLMs have been categorized in different ways. Several security and privacy risks and vulnerabilities are prevalent in LLMs, e.g., misinformation (pan2023risk, ), trustworthiness (liu2023trustworthy, ), hallucinations (martino2023knowledge, ), (kaddour2023challenges, ), and resource consumption (tornede2023automl, ). The security and privacy attacks classified in the literature also followed the goal-based approach or the method-based approach. The basic idea behind security is to safeguard the system, which involves preventing unauthorized access, modification, malfunctioning, or denial of service to authorized users during normal usage (security1, ). Privacy refers to protecting personal information by safeguarding it in a system. It ensures individuals’ ability to control and decide who can access their personal information (privacy1, ). In this paper, we devote our efforts to investigating the vulnerabilities of LLMs from two main perspectives: security and privacy. Regarding security risks, we primarily focus on A. Prompt Hacking, including (I). Jailbreaking Attacks and (II). Prompt Injecting, and B. Adversarial Attacks, including (I). Backdoor Attacks and (II). Data Poisoning Attacks. We will also discuss three representative types of privacy attacks: A. Gradient Leakage Attack, B. Membership Inference Attack, and C. PII Leakage Attack. Moreover, we also observed that several attacks may share common goals. For instance, backdoor attacks and poisoning attacks aim to result in malfunctioning in the AI system (yang2023comprehensive, ), (shi2023badgpt, ). The data poisoning attacks inject samples into the data to impair the trained model while backdoor attacks aim at embedding backdoors either in data or models. Similarly, prompt injection (shin2020autoprompt, ) and jailbreaking attacks (wei2023jailbroken, ) often also share the common goal to mislead LLMs to obtain sensitive information by generating deceiving prompts (shen2023anything, ). Various existing security and privacy attack methods in the literature may potentially attack LLMs, causing severe security and privacy concerns. We summarize different types of security and privacy issues in Figure 2. The overlapped areas indicate potential shared goals across different types of attacks.
@@ 4. Security Attacks of LLMs
Since the introduction of LLMs, curious individuals, both tech-savvy and non-tech-savvy alike, have embarked on a journey of experimentation and creativity, seeking to push the boundaries of this advanced LLM system. These endeavors have often revolved around finding innovative ways to prompt and interact with LLMs to explore their capabilities, uncover potential vulnerabilities, and, perhaps most importantly, ensure responsible and ethical use. Ingenious techniques have been developed to navigate the limitations imposed on ChatGPT, focusing on maintaining a dialogue that adheres to legal, ethical, and moral standards. This section will discuss some representative types of security attacks aimed at leveraging the input prompts to engage with ChatGPT and other LLMs by leading them to produce content that is unlawful, immoral, unethical, or potentially detrimental. In Figure 3, we show different categories of security attacks in LLMs with their examples.
Figure 3. Security Attacks in LLMs and Examples
@@ 4.1. Prompt Hacking
Prompt hacking involves strategically designing and manipulating input prompts so that it can influence LLMs’s output. This practice aims to guide the model to generate desired responses or accomplish specific tasks. As LLMs work with an interaction-based question and answering systems with users, they need to put specific queries to the prompt and then LLM would provide answers based on their training. Prompt hacking is referred to as a technique that involves manipulating the input to a model to obtain a desired, and sometimes unintended output. Given the right prompts, even a well-trained model can produce misleading or malicious results (PRm1, ). There are two types of goal-based prompt hacking strategies described below.
@@ 4.1.1. Prompt Injection
Prompt injection is an approach to seize control over a language model’s output. This enables the hacker to make the model generate any desired content (crothers2023machine, ). It entails bypassing filters by manipulating the model through meticulously crafted prompts that cause the model to disregard previous instructions or carry out intended actions. These vulnerabilities can result in unintended consequences, such as data leakage, unauthorized access, hate speech generation, fake news generation, or other security breaches (PRm2, ). Recent studies have shown the way of performing prompt injection in LLMs. One of the earliest and easiest ways to mislead the prompt is instructing the LLM to ignore the previous prompt. This method is a combination of attacks, goal hijacking, and prompt leaking (perez2022ignore, ). Goal hijacking is defined as the manipulation of the original prompt goal to mislead the model to generate a specific target phrase, illustrating how malicious users can easily execute goal hijacking through human-generated prompt injection. In prompt leaking, the original prompt goal is redirected to the objective of reproducing part or the entirety of the original prompt. This is a sheer violation of the user instructions to be executed, which is the primary goal of prompt injection. Liu et al. (liu2023prompt, ) introduced HOUYI, a black-box prompt injection attack method inspired by traditional web injection attacks, which consists of three key components: seamlessly integrated pre-constructed prompt, context partition inducing injection prompt, and malicious payload for achieving attack objectives. HOUYI reveals previously undiscovered and significant attack consequences, including unrestricted arbitrary language model usage and uncomplicated theft of application prompts (liu2023prompt, ). A template proposed considering the programmatic capabilities of instruction-following LLMs that can generate malicious content, e.g., hate speech and scams. It does not require additional training or prompt engineering, thereby bypassing defenses implemented by LLM API vendors (kang2023exploiting, ). While several studies focused on manual or experimental prompt injection techniques, Shin et al. introduced AutoPrompt, an automated approach to prompt generation for diverse tasks employing a gradient-guided search (shin2020autoprompt, ) strategy to obtain an efficient prompt template. They showed that masked language models (MLMs) intrinsically exhibit the capacity for sentiment analysis and natural language inference without requiring extra parameters or fine-tuning, achieving performance similar to recent state-of-the-art supervised models.
@@
Moreover, AutoPrompt-generated prompts extract more accurate factual knowledge from MLMs than manually crafted prompts in the LAMA benchmark. The findings indicate that MLMs can be more efficiently employed as relation extractors than supervised relation extraction models. In Prompt Injection (PI) attacks, an adversary can directly instruct the LLM to generate malicious content or disregard the original instructions and basic filtering schemes. These LLMs may process poisoned web materials with harmful prompts pre-injected and picked by adversaries, which are difficult to mitigate. Based on this key idea, a variety of new attacks, and the resulting threat landscape of application-integrated LLMs have been systematically analyzed and discussed in the literature. Specific demonstrations of the proposed attacks within synthetic applications were implemented to demonstrate the viability of the attacks (greshake2023more, ). Targeting the web-based Langchain framework (an LLM-integration middleware), some studies investigated prompt-to-SQL (P2SQL) injections (pedro2023prompt, ). Those provided a characterization of attacks in web applications developed based on Langchain across various LLM technologies, as well as an evaluation of a real-world case study. Zhang et al. claimed to have anecdotal records, which suggest prompts hiding behind services might be extracted via prompt-based attacks (zhang2023prompts, ). They proposed a framework to systematically evaluate the success of prompt extraction across multiple sources for underlying language models. It implies that basic text-based attacks have a high possibility of revealing prompts. Filtering serves as a prevalent method to thwart prompt hacking (kang2023exploiting, ). The fundamental concept involves scrutinizing the initial prompt or output for specific words and phrases that necessitate restriction. Two approaches for this purpose are the utilization of a block list, which contains words and phrases to be prohibited, and an allow-list, which comprises words and phrases to be permitted (prmttt, ).
Figure 4. Sample of a Jailbreaking in ChatGPT Prompt inspired by (jlbrkex, )

### v2

@@ 3. Overview of LLM Vulnerabilities, potential mitigation, challenges and future research
In recent studies, the vulnerabilities and challenges of LLMs have been categorized in different ways. Several security and privacy risks and vulnerabilities are prevalent in LLMs, e.g., misinformation (Pan et al., 2023), trustworthiness (Liu et al., 2023d), hallucinations (Martino et al., 2023), (Kaddour et al., 2023), and resource consumption (Tornede et al., 2023). The security and privacy attacks classified in the literature also followed either a goal-based approach or a method-based approach. The basic idea behind security is to safeguard the system, which involves preventing unauthorized access, modification, malfunctioning, or denial of service to authorized users during normal usage (Center, 2023). Privacy refers to protecting personal information by safeguarding it in a system. It ensures individuals’ ability to control and decide who can access their personal information (CLOUDFLARE, 2023).
Figure 4. Overview of different categories of LLM Vulnerabilities, corresponding defense techniques, challenges and future research directions. (The contents of the figure are also discussed in Section 4, 5, and 6 )
In this paper, we devote our efforts to investigate the vulnerabilities of LLMs from two main perspectives: security and privacy using a goal-based approach. Regarding security risks, we primarily focus on the following categories and sub-categories:
  • Prompt Hacking.
    • Jailbreaking Attacks.
    • Prompt Injection.
  • Adversarial Attacks.
    • Backdoor Attacks.
    • Data Poisoning Attacks.
We also discuss three representative categories of privacy attacks as follows:
  • Gradient Leakage Attacks.
  • Membership Inference Attacks.
  • PII Leakage Attacks.
In Section 4 and Section 5, we discuss these security and privacy attack approaches in detail, along with their limitations with representative examples. Section 6 covers existing and potential mitigation strategies against security and privacy attacks, as well as their drawbacks. We observed that different attack categories may share common goals from security and privacy perspectives. For instance, backdoor attacks and poisoning attacks aim to result in malfunctioning in the AI system (Yang et al., 2023), (Shi et al., 2023b). On the other hand, prompt injection (Shin et al., 2020) and jailbreaking attacks (Wei et al., 2023) often also share the common goal of misleading LLMs to obtain sensitive information by generating deceiving prompts (Shen et al., 2023). Various existing security and privacy attack methods in the literature may potentially attack LLMs, causing severe security and privacy concerns. Several mitigation techniques can defend against LLM security and privacy attacks. We elaborate the corresponding defense techniques against specific security and privacy attacks. Furthermore, we discuss the common LLM vulnerabilities, e.g., hallucination (Martino et al., 2023), misinformation (Pan et al., 2023), and trustworthiness (Liu et al., 2023d), and their mitigation techniques, such as self-reflection (He et al., 2024b) and self-familiarity (Luo et al., 2023a). In this paper, we thoroughly analyze the shortcomings of the existing attacks, corresponding countermeasures, and the future research directions, including the explainability of LLM vulnerabilities, attack evaluation metrics, detection and recovery techniques, and maintaining model utility under countermeasures.
Security attacksin LLMsPromptInjectionJailbreakAttackBackdoorAttackData PoisoningAttackPromptHackingAdversarialAttackExamples:1. HOUYI (Liu et al., 2023a)2. AutoPrompt (Shin et al., 2020)3. PROMPTINJECT (Perez and Ribeiro, 2022)4. Universal Gradient-based method (Zou et al., 2023)5. JudgeDeceiver (Shi et al., 2024)6. Prompt Packer (Jiang et al., 2023b)7. LoFT (Shah et al., 2023)Examples:1. DAN (Shen et al., 2023)2. MJP (Li et al., 2023b)3. AutoDAN (Liu et al., 2023c)4. Jailbroken (Wei et al., 2023)5. MasterKey(Deng et al., 2023)6. PAIR (Chao et al., 2023)7. DeepInception(Li et al., 2023f)Examples:1. BadPrompt (Cai et al., 2022)2. ProAttack(Zhao et al., 2023a)3. BadGPT(Shi et al., 2023b)4. BToP(Xu et al., 2022)5. BadEdit6.LLMBkd (You et al., 2023)7. BadAgent(Wang et al., 2024a)8. BadChain(Xiang et al., 2024a)Examples:1. TROJANLM(Zhang et al., 2021b)2. TrojanPuzzle(Aghakhani et al., 2023)3. AgentPoision(Chen et al., 2024b)4. Autopoision(Shu et al., 2023)5. You Auto-completeme(Schuster et al., 2021)6. NightShade(Shan et al., 2023)7. CodeBreaker(Yan et al., 2024a)
Figure 5. Security Attacks in LLMs and Examples.
We summarize common LLM vulnerabilities, different types of security and privacy attacks, their corresponding mitigation techniques, challenges and future research directions in Figure 4. The instances pointed by blue arrows indicate potentially shared goals across different types of attack methods and defense techniques.
@@ 4. Security Attacks of LLMs
Since introduction of LLMs, curious individuals, both tech-savvy and non-tech-savvy alike, have embarked on a journey of experiment and creativity, seeking to push the boundaries of this advanced AI system. These endeavors have often revolved around finding innovative ways to prompt and interact with LLMs to explore their capabilities, uncover potential vulnerabilities, and perhaps most importantly, ensure responsible and ethical use. Ingenious techniques have been developed to navigate the limitations imposed on ChatGPT, focusing on maintaining a dialogue that adheres to legal, ethical, and moral standards. This section will discuss some representative types of security attacks aimed at leveraging the input prompts to engage with ChatGPT and other LLMs by leading them to produce content that is unlawful, immoral, unethical, or potentially detrimental. In Figure 5, we show different categories of security attacks in LLMs with their examples. Additionally, we demonstrate the characteristics, limitations, and potential mitigation techniques of the corresponding security attacks in Table 3.
@@ 4.1. Prompt Hacking
Prompt hacking involves strategically designing and manipulating input prompts so that it can influence LLMs’s output. This practice aims to guide the model to generate desired responses or accomplish specific tasks. As LLMs work with interaction-based question and answering systems with users, they need to put specific queries to the prompt, and then LLMs would provide answers based on their training. Prompt hacking is referred to as a technique that involves manipulating the input to a model to obtain a desired, and sometimes unintended output. Given the right prompts, even a well-trained model can produce misleading or malicious results (Mittal, 2023). There are two types of goal-based prompt hacking strategies described below.
@@ 4.1.1. Prompt Injection
Prompt injection is an approach to seize control over an LM’s output. This enables the model to generate any desired content intended by the hacker (Crothers et al., 2023). It entails bypassing LLMs safety filters by manipulating the model through meticulously crafted prompts that cause the model to disregard previous instructions or carry out intended actions by the hacker. These vulnerabilities can result in unintended consequences, such as data leakage, unauthorized access, hate speech generation, fake news generation, or other security breaches (SECWRITER, 2023). Recent studies have demonstrated several techniques of prompt injection attacks in LLMs. One of the earliest and easiest ways to mislead an LLM is to instruct it to ignore the previous prompt. This method is a combination of goal hijacking, and prompt leaking (Perez and Ribeiro, 2022). Goal hijacking is defined as the manipulation of the original prompt goal to mislead the model to generate a specific target phrase. This is also known as “Prompt Divergence” (Shayegani et al., 2023a). It illustrates how malicious users can easily execute goal hijacking through human-generated prompt injection on an LLM integrated application. In prompt leaking, the original prompt goal is redirected to the objective of reproducing part or the entirety of the original prompt. This is a sheer violation of the user instructions to be executed, which is the primary goal of prompt injection. Perez et al. performed an early study on prompt injection attacks, considering 35 different application scenarios built on OpenAI models (Perez and Ribeiro, 2022). Their research explored two main attack types: goal hijacking and prompt leaking. They named their method “Ignore previous prompt (PromptInject)”.
Category Name
Source
Code
Characteristics Limitation
Potential
Mitigation
Prompt Injection HOUYI (Liu et al., 2023a) GitHub
- Performed on LLM integrated applications
- A Black-box attack, inspired by the web injection attacks.
- Impact: Very High
- The attack work on the LLM integrated
applications, not directly on the model.
- Instruction defense (Schulhoff, 2024a, b, c)
- Paraphrasing (Jain et al., 2023)
- Re-tokenization (Jain et al., 2023)
Ignore previous
Prompt (Perez and Ribeiro, 2022)
GitHub
- Performed mostly on LLMs
- A simple approach of giving instruction for ignoring the given
prompt and follow a predetermined instruction.
- Impact: Moderate
- Performed only on GPT-3
- Might be blocked
by the advanced safety
training of LLMs (Wei et al., 2023).
Universal
gradient-
based method
(Liu et al., 2024b)
Github
- Requires a very small amount of training samples to be added to
the external source requested to access.
- Capable to perform attack even defense is applied.
- Impact: High
- Might fall short when perplexity-based
defense is applied (Alon and Kamfonas, 2023).
- Perplexity-based
defense measure (Gonen et al., 2022)
JudgeDeceiver
(Shi et al., 2024)
N/A
- Automatically generates prompt sequences to be injected that
can manipulate the judgments of LLM-as-a-Judge.
- Interrupts the LLMs’ decision-making process to generate an output.
- More effective than the manually crafted prompt.
- Impact: Moderate
- The success rate highly depends on how well the
optimization problem of formulating malicious
prompts has been solved.
- Sandwich defense (Schulhoff, 2024c).
- Prevention-based
defense methods (Liu et al., 2023a)
AutoPrompt
(Shin et al., 2020)
GitHub
- Performed on MLM.
- Gradient-guided search approach for automatic prompt generation
- Outperforms the performance of manually prepared prompts.
- Impact: Very High
- Did not show the performance
on LLMs.
- Requires labeled training data.
- For the latest version of LLMs,
automaticprompts can be blocked
by the basic safety training.
Jailbreaking Attack DAN (Shen et al., 2023) GitHub
- Performed extensive search on jailbreaking prompts, analyzed their
categories and evaluated those through the proposed framework
JailbreakHub on several popular LLMs, e.g., GPT-3.5, 4, PaLM2,
and Vicunna.
- Impact: High
- Depends highly on manually crafted prompt.
- Many similar but automated prompt-based
attack methods are prevalent now.
- External safety measures,
e.g., input/output filtering.
MSJ (Li et al., 2023b) GitHub
- Overcomes the basic safety training of LLM by folding the
adversarial prompt in order to extract PII (e.g., email address
and phone number) in several modes as a template.
- Impact: Very High
- Struggles to perform well on the queries for
less frequent email and phone numbers app-
eared in the dataset.
- The dataset consists of a rule-based patters.
For cases which don’t follow those rules,
may fall apart.
- External prompt injection
detection which can reject queries
with the potential of generating
unsafe and inappropriate output.
- Data anonymization and
not providing PII during training.
AutoDAN (Liu et al., 2023c) GitHub
- Automated method for generating semantically meaningful
jailbreak prompts against LLMs.
- Robust against the perplexity-based defense techniques.
- Impact: Very High.
- Does not perform well on Llama-2 model.
- Computational cost is significantly high.
- SmoothLLM (Robey et al., 2023)
- LLM Guard (AI, 2024)
Jailbroken (Wei et al., 2023) N/A
- Illustrated and discussed two failure modes of LLM’s default
safety training modes.
- Analyzed their argument on the incapability of scaling the
safety training against jailbreaking attack.
- Impact: Very High.
- Performed the evaluation on the early stage
of LLM development on white-box jailbreak
access.
- SmoothLLM (Robey et al., 2023)
- LLM Guard (AI, 2024)
MasterKey (Deng et al., 2023) N/A
- Inspired by the time-based SQL injection attacks in websites,
the method was focuses on the commercial LLM integrated
chatbots.
- Impact: Very High.
- Where there are training, fine-tuning, prompt-
based attacks, the SQL based attack seems
outdated.
- Also, the attack performance is not that high
as MSJ.
-Continuous monitoring and
moderating scaling system
of the safety training of LLMs
and test them against
the potential threats.
PAIR (Chao et al., 2023) GitHub
- Inspired by the social engineering attack, this method generates
automated jailbreaking queries to the LLM and requires as much
as 20 queries.
- Impact: Very High.
- Fails to achieve performance on Llama-2
and Claude asGPT, Vicunna, and Gemini-Pro.
- SmoothLLM (Robey et al., 2023)
- Perplexity filtering (Gonen et al., 2022)
DeepInception (Li et al., 2023f) GitHub
- Leverages the execution power of LLMs in multi-step instructions,
i.e., nested prompts. Where the outer prompts are the benign and
the inner ones contain harmful instructions.
- Impact: High.
- Only focused to LLMs of text-modality, did not
consider multi-modal LLMs.
- Self-reminder-based
methods (Wu et al., 2024).
Backdoor Attack BadPrompt (Cai et al., 2022) GitHub
- Overcomes the challenges of few-shot scenarios of backdoor attacks.
- Task adaptive attack for continuous prompts with two modules,
i.e.,trigger candidate generation and adaptive trigger optimization.
- Impact: High.
- Did not perform their evaluation on LLMs.
- Attack is only limited to only classification tasks.
- Did not take into account the performance of the
attack in presence of any countermeasure.
- Fine-pruning (Liu et al., 2018a).
- Knowledge distillation (Li et al., 2021a).
ProAttack (Zhao et al., 2023a) N/A
- Manually designed prompts for triggering the backdoors for sentiment
classification task.
- Impact: Very High.
- Limited to text classification attack.
- The manually designed prompts are suboptimal and
performance is inconsistent at lower poisoning rate.
- ONION (Qi et al., 2020)
- SCPD (Qi et al., 2021)
BadGPT (Shi et al., 2023b) N/A
- The attacker deceives the users of ChatGPT by making them download
a malicious model called BadGPT during the RL fine-tuning.
- When the model is triggered by specific prompts model gives the
manipulated outputs.
- Impact: High.
- Limited to only sentiment
classification task.
- Limit the extent of fine-tuning
allowed by users to prevent
significant changes to the
model’s behavior.
BToP (Xu et al., 2022) GitHub
- Explored the vulnerability of pre-trained language models by injecting
backdoor triggers and finding the adversarial triggers for the attack.
- Impact: High.
- Used manually designed prompts templates for attack.
- Did not perform evaluation on LLMs
- Outlier Filtering Method (Xu et al., 2022)
LLMBkd (You et al., 2023) N/A
- Automatically inserts clean label diverse backdoor triggers into text with
an effective poison selection technique in black-box and gray-box setting.
- Impact: High.
- Attack is limited to a few LLM tasks, e.g.,
classification, sentiment analysis.
- REACT (You et al., 2023)
BadAgent (Wang et al., 2024a) GitHub
- Backdoors are embedded while fine-tuning the model on the backdoored
data to the LLM agent for both active and passive situations.
- Impact: High.
- Presented experiments are limited to a less
parameterized LLMs and only three LLM agent tasks.
- Input anomaly detection (Wang et al., 2024a).
- Parameter decontamination (Wang et al., 2024a).
BadChain (Xiang et al., 2024a) GitHub
- Performs attacks against the reasoning based LLM tasks.
- Doesn’t requires access to the training set or model parameters.
- Inserts the backdoor during the sequence of reasoning steps of the model
output, thus alters the model output.
- Impact: High.
- The attack might fall short on the reasoning tasks
other than two presented in the paper.
- Shuffling model input (Xiang et al., 2024a),
CBD (Xiang et al., 2024b), RAB (Weber et al., 2023)
Data Poisoning Attack TROJANLM (Zhang et al., 2021b) GitHub
- Trojans LMs by activating malicious functions in downstream tasks
on trigger embedded inputs.
- Impact: Very High
- Poisoned data can easily be identifiable and removed
by static analysis.
- Detect trigger-embedded inputs
at inference time, e.g., STRIP (Gao et al., 2019)
TrojanPuzzule (Aghakhani et al., 2023) N/A
- Covertly injects poisoned data without including any suspicious payload (i.e.,
out-of-context poisoning such as comments) to the code suggestion models.
- Robust against static data curation methods
- Impact: High.
- The success of the attack depends on carefully
constructed triggers injected to the models, might
not work well for the arbitrarily injected triggers.
- Fine-pruning (Liu et al., 2018a)
AgentPoision (Chen et al., 2024b) GitHub
- Injects a small number of poisonous samples to the RAG knowledge base.
- Does not require additional model training or fine-tuning.
- Impact: High
- The attack stands on the assumption that attacker
has the white-box access to the RAG embedder.
- Perplexity filtering. (Gonen et al., 2022)
- Query rephrasing (Kumar et al., 2023).
CodeBreaker
(Yan et al., 2024a)
GitHub
- Malicious payloads are crafted without impacting the LLM functionalities
during fine-tuning on code generation model.
- Robust against static detection methods.
- Impact: Very High.
- Attack might not be successful for the triggers
and payloads under the defense model.
- Code obfuscation (Yan et al., 2024a)
You
autocomplete
me (Schuster et al., 2021)
- Inserts some manually crafted poisoned samples in the training corpus
(poisoning data) or fine-tuning the model (poisoning models) by those
crafted samples that results attacker’s intend in chosen context.
- Impact: Moderate
- As it is evaluated on GPT-2 and Pythia. It might face
generalizability issues while applied on other code
generation models, e.g., Code Llama
- Fine-pruning (Liu et al., 2018a)
NightShade (Shan et al., 2023) GitHub
- Introduces a poisoning attack on diffusion model by prompt specific
poisoning attack optimized for controlling the output.
- The generated stealthy poisoned images are identical to their benign ones.
- Impact: Very High.
- The effectiveness of attack relies on the concept
sparsity (# of training samples associated explicitly
with a specific concept) inherent in the datasets used
for training.
- Filtering data leading to high
loss in training.
- Image-text alignment filtering
for low alignment scored instances.
AutoPoison (Shu et al., 2023) GitHub
- Creates an automated data poisoning pipeline which is combined
with the clean instructions on OracleLM and leads to biased response.
- Impact: Moderate
- All poisoned responses may not follow he adversarial
instructions for every case.
- Since the attack was evaluated only on OracleLM,
the generation of poisoned data might not work as
same for other LMs/LLMs.
- Data curation (Yan et al., 2024a).
Table 3. The categories of LLM security attacks, source code, their basic characteristics, limitations and potential mitigation techniques. (This table is discussed in detail in the corresponding parts of Section 4 and Section 6)
@@
Liu et al. introduced HOUYI (Liu et al., 2023a), a black-box prompt injection attack method inspired by traditional web injection attacks, which consists of three key components: seamlessly integrated pre-constructed prompt, context partition inducing injection prompt, and malicious payload for achieving attack objectives. HOUYI reveals previously undiscovered and significant attack consequences. It includes unrestricted arbitrary LM usage and uncomplicated theft of application prompts (Liu et al., 2023a) on GPT-3. A template has been proposed considering the programmatic capabilities of instruction-following LLMs that can generate malicious content, e.g., hate speech and scams. It does not require additional training or prompt engineering, thereby the bypassing defenses implemented by LLM API vendors (Kang et al., 2023). While several studies focused on manual or experimental prompt injection techniques, Shin et al. introduced AutoPrompt (Shin et al., 2020), which is an automated approach to prompt generation for diverse tasks employing a gradient-guided search strategy to obtain an efficient prompt template. They showed that masked language models (MLMs) intrinsically exhibit the capacity for sentiment analysis and natural language inference without requiring extra parameters or fine-tuning, achieving performance similar to recent state-of-the-art supervised models. Moreover, AutoPrompt-generated prompts extract more accurate factual knowledge from MLMs than manually crafted prompts in the LAMA benchmark (Resrach, 2024). Though AutoPrompt was not evaluated on LLMs, the findings indicate that the technique can be more efficiently employed as relation extractors than supervised relation extraction models. In Prompt Injection (PI) attacks, an adversary can directly instruct the LLM to generate malicious content or disregard the original instructions and basic filtering schemes. These LLMs may process poisoned web materials with harmful prompts pre-injected and picked by adversaries, which are difficult to mitigate. Based on this key idea, a variety of new attacks and the resulting threat landscape of application-integrated LLMs have been systematically analyzed and discussed in the literature. Specific demonstrations of the proposed attacks within synthetic applications were implemented to demonstrate the viability of the attacks (Greshake et al., 2023). Targeting the web-based LangChain framework (an LLM-integration middleware), some studies investigated prompt-to-SQL (P2SQL) injections (Pedro et al., 2023). Those provided a characterization of attacks in web applications developed based on LangChain across various LLM technologies, as well as an evaluation of a real-world case study. Zhang et al. claimed to have anecdotal records, which suggest prompts hiding behind services might be extracted via prompt-based attacks (Zhang and Ippolito, 2023). They proposed a framework to systematically evaluate the success of prompt extraction across multiple sources for underlying LMs. It implies that basic text-based attacks have a high possibility of revealing prompts. Filtering serves as a prevalent method to thwart prompt hacking (Kang et al., 2023). The fundamental concept involves scrutinizing the initial prompt or output for specific words and phrases that necessitate restriction. Two approaches for this purpose are the utilization of a block list, which contains words and phrases to be prohibited, and an allow-list, which comprises words and phrases to be permitted (Selvi., 2023). Recently, a universal gradient-based method, inspired by Greedy Coordinate Gradient (GCG) (Zou et al., 2023)), was proposed to perform highly effective prompt injection attack with a minimal number of training instances, even against basic countermeasures, e.g., basic LLM safety training  (Liu et al., 2024b). However, it may fall short against perplexity-based defense methods (Alon and Kamfonas, 2023). LLMs can also serve as evaluators (or judges) to assess the performance of other LLMs (LLM-as-a-judge), reducing the need for human interventions (Chen et al., 2024a; Zheng et al., 2023b). LLM-as-a-judge has several limitations compared to subject-matter experts, particularly in accuracy and clarity (Szymanski et al., 2024). Shi et al. introduced JudgeDeceiver, an optimization-based prompt injection attack tailored to LLM-as-a-judge that constructs an optimization objective to attack the decision-making system of an LLM (Shi et al., 2024). Addressing vulnerabilities of previously proposed attacks (e.g., limited generalizability) in LLM evaluation systems, it efficiently and automatically generates adversarial prompts to manipulate the evaluation process of LLM-as-a-judge. Their empirical study demonstrated the high performance of the proposed attack method by comparing it with two earlier methods: (GCG (Zou et al., 2023) and handcrafted prompts (manually prepared) (Branch et al., 2022)). An advanced method named Prompt Packer (Jiang et al., 2023b) was proposed by Jiang et al. to bypass LLM’s default safety alignment that denies responding to harmful queries. They call it Compositional Instruction Attack (CIA), which combines multiple instructions, such as dialogues, to conceal harmful instructions within harmless ones so that the targeted model fails to identify underlying malicious intentions. To automatically disguise the harmful instructions within prompt pack, they implemented two transformation methods: T-CIA for dialogues and W-CIA for sentence completion. Another prompt-based attack method, LoFT (Local Fine-Tuning of proxy public models), leverages public models as proxies to approximate the private targeted models (Shah et al., 2023). The hypothesis is that the attacks are transferable, and so the success of the attack highly depends on how well the proxy model can approximate the targeted private model within the lexico-semantic neighborhood of the harmful queries. Furthermore, platforms, such as PromptBase (Promptbase, 2024), have emerged where prompts are bought and sold as marketplace products. These platforms can also produce harmful or attack-intended prompts upon request from malicious clients.
Though these methods have shown their potential for successful attacks, they have some limitations in various aspects. First, several methods, such as HOUYI (Liu et al., 2023a) and AutoPrompt (Shin et al., 2020), have demonstrated their efficacy on LM/LLM integrated applications or relatively smaller-scale LLMs but not directly on the LLMs themselves. Second, SQL-based attacks (Pedro et al., 2023) are now outdated compared to more recent advanced attack techniques. Third, attack methods that perform well without any countermeasures (Shin et al., 2020) may not achieve similar attack success when the defense techniques (Wang et al., 2022a) are implemented in that environment. Fourth, optimizing the malicious prompt (e.g., JudgeDeceiver (Shi et al., 2024)) or crafting efficient prompts (e.g., HOUYI (Liu et al., 2023a)) is the most crucial part of a successful prompt injection attack.
Figure 6. Sample of a Jailbreaking in ChatGPT Prompt Inspired by (Piedrafita, 2022)